OpenStudy (anonymous):

A radioactive substance decays at a continuous rate of 2% per year. Suppose that 5 grams are present as of now. a) how many grams will be present after 10 years? b) find the half-life.

OpenStudy (anonymous):

\[ a(t)=a_0(100\%-2\%)^t \]

OpenStudy (anonymous):

First of all, when they say decrease by \(2\%\), that tells you that the new percent it will be is \(100\%-2\%\).

OpenStudy (anonymous):

ok so 98%

OpenStudy (anonymous):

Which evaluates to \(0.98\).

OpenStudy (anonymous):

When \(t=0\), then it will evaluate to \(1\). So that means \(a_0\) must be the initial amount.

OpenStudy (anonymous):

So I guess, we can plug that in: \[ a(t) = 5(0.98)^t \]

OpenStudy (anonymous):

And they want you to find \(a(10)\) for the first part.

OpenStudy (anonymous):

For the second part, with the half life we want to set \(a(t) = a_0/2\), that is, half the initial amount. Then we can solve for \(t\).

OpenStudy (anonymous):

So that leaves us with: \[ 2.5=5(9.8)^t \]Solving for \(t\).

OpenStudy (anonymous):

Am I going to fast, or do you think you got it now?

OpenStudy (anonymous):

so for the first part would i solve it by calculating \[5(0.98)^{10}\] ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

and the second part the same thing?

OpenStudy (anonymous):

Second part is solving for \(t\).

OpenStudy (anonymous):

can you help explain how i solve for t for the second part?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Logarithms.

OpenStudy (anonymous):

this is the part i get confused on...

OpenStudy (anonymous):

Okay, first: \[ 2.5=5(0.98)^t \]What is the first thing you want to do?

OpenStudy (anonymous):

divide from both sides?

OpenStudy (anonymous):

Yes, divide by \(5\) first.

OpenStudy (anonymous):

.5

OpenStudy (anonymous):

Okay, so \[ \frac{1}{2}=(0.98)^t \]

OpenStudy (anonymous):

Now do natural log of both sides: \[ \ln(1/2) = \ln(0.98^t) \]

OpenStudy (anonymous):

Do you know how to simplify this?

OpenStudy (anonymous):

ln(1/2) divided by ln(0.98)?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

34.30961849?

OpenStudy (anonymous):

Yes, that approximate.

OpenStudy (anonymous):

whats that rounded to the nearest hundredth?

OpenStudy (anonymous):